### v2

# A class of forward-backward regularizations of the Perona-Malik equation with variable exponent

Abstract
This paper investigates a novel class of regularizations of the Perona-Malik equation with variable exponents, of forward-backward parabolic type, which possess a variational structure and have potential applications in image processing. The existence of Young measure solutions to the Neumann initial-boundary value problem for the proposed equation is established via Sobolev approximation and the vanishing viscosity limit. The proofs rely on Rothe’s method, variational principles, and Young measure theory. The theoretical results confirm numerical observations concerning the generic behavior of solutions with suitably chosen variable exponents.
keywords:
Forward-backward parabolic equations , Pseudo-parabolic regularization , Nonstandard growth condition , Young measure solutions
journal: arXiv
1 Introduction
In this paper, we consider the following forward-backward parabolic problem with nonstandard growth condition:
$\displaystyle\frac{\partial u}{\partial t}=\operatorname{div}\left(\left(\frac{1}{1+\left|\nabla u\right|^{2}}+\delta\left|\nabla u\right|^{p(x)-2}\right)\nabla u\right),$ $(x,t)\in Q_{T}$, (1)
$\displaystyle\frac{\partial u}{\partial\vec{n}}=0,$ $(x,t)\in\partial\Omega\times(0,T)$, (2)
$u(x,0)=f(x),$ $x\in\Omega$, (3)
where $\Omega\subset\mathbb{R}^{N}$ ($N\geq 2$) is a bounded domain with smooth boundary $\partial\Omega$, $0<T<\infty$, $Q_{T}=\Omega\times(0,T)$, $\vec{n}$ denotes the unit outward normal to $\partial\Omega$, and $\delta>0$. The variable exponent function $p$ is globally $\log$-Hölder continuous and satisfies
$2\leq p^{-}=\operatorname*{ess\,inf}_{z\in\Omega}p(z)\leq p(x)\leq p^{+}=\operatorname*{ess\,sup}_{z\in\Omega}p(z)<\infty.$ (4)
When $p>1$ is a constant, equation (1) arises in [Guidotti20131416] as a model for image denoising, where $u$ denotes the restored image, $f$ the noisy observation, and the function values quantify the gray levels. For sufficiently small $\delta$, this equation is of forward-backward parabolic type and admits no classical strong or distributional solutions. In [Kinderlehrer1992001], Young measure theory was employed to study the Dirichlet initial-boundary value problem (IBVP) for equation (1) with $p=2$, where the authors proved the existence of Young measure solutions using Rothe’s method and variational techniques combined with relaxation theory [Dacorogna2008]. Demoulini [Demoulini1996376] further investigated the qualitative properties of these solutions, including uniqueness, the comparison principle, and long-time behavior, which were later extended from the Dirichlet to the periodic and Neumann settings in [GUIDOTTI20123226]. The well-posedness of the Dirichlet IBVP for equation (1) was generalized to the case $1<p\leq 2$ by adopting an approach similar to that in [Demoulini1996376]. For the case $p=1$, equation (1) becomes singular, and Young measure solutions are considered in $L^{\infty}(0,T;BV(\Omega))$; see [Yin2003659]. In fact, equation (1) with constant exponent $p>1$ serves as a forward-backward regularization of the well-known Perona-Malik equation
$u_{t}=\operatorname{div}\left(\frac{\nabla u}{1+\left|\nabla u\right|^{2}}\right),$ (5)
which is parabolic if $\left|\nabla u\right|<1$ and backward-parabolic if $\left|\nabla u\right|>1$. Originally introduced for image reconstruction, the Perona-Malik equation is designed to reduce noise while preserving important image features, particularly edges and contours [Perona1990629]. In practice, however, the Perona-Malik equation suffers from the drawback that it may result in staircase artifacts in the restored image. This issue inherently stems from the ill-posedness of (5), whose instabilities during the evolution manifest as the so-called staircasing phenomenon in numerical discretizations [Guidotti2015131]; see Figure 1. From a dynamical perspective, this phenomenon is replaced by a milder, microstructured ramping in the $p$-Laplacian regularized Perona-Malik equation, characterized by alternating gradients of finite amplitude; see Figure 2.
Forward-backward parabolic equations arise in various applications, including geophysics, biology, and the modeling of physical phenomena. A typical form can be written as
$u_{t}=\operatorname{div}\vec{\varPhi}(\nabla u),$ (6)
where the heat flux $\vec{\varPhi}(\xi)=\nabla\varPsi(\xi)$ for some nonconvex potential $\varPsi\in C^{1}(\mathbb{R}^{N})$. In this case, the monotonicity condition $(\vec{\varPhi}(\xi)-\vec{\varPhi}(\eta))\cdot(\xi-\eta)\geq 0$ is violated for some $\xi,\eta\in\mathbb{R}^{N}$. The one-dimensional version of equation (6) reads
$u_{t}=\left[\varPhi(u_{x})\right]_{x},$ (7)
and appears independently as a model for heat and mass transfer in a stably stratified turbulent shear flow [Barenblatt1993341]. Setting $w=u_{x}$, formal differentiation of (7) with respect to $x$ leads to another type of forward-backward parabolic equation
$w_{t}=\left[\varPhi(w)\right]_{xx},$ (8)
which arises in the theory of phase transitions [Brokate2012]. In [Hollig1983299], the Neumann IBVP for (7) is shown to admit infinitely many weak $L^{2}$-solutions when $\varPhi$ is a nonmonotonic piecewise linear function satisfying $s\varPhi(s)\geq Cs^{2}$ for some $C>0$. This condition on $\varPhi$ was later relaxed in [ZHANG2006322] by assuming higher regularity of the initial data. In [Zhang2006171], the existence of infinitely many weak $W^{1,\infty}$-solutions to the one-dimensional Perona-Malik equation under homogeneous Neumann boundary conditions was established. As shown in [GURTIN1996178, Kichenassamy19971328], no global $C^{1}$ solution exists for sufficiently large $\left|f^{\prime}\right|$, which suggests that even the local existence of solutions to the IBVP for equations (6) or (7) remains a nontrivial problem. Concerning the uniqueness of forward-backward parabolic problems, the situation becomes more subtle (see [LAIR1985311]). For more theoretical results related to the Perona-Malik equation, the reader is referred to [BOTTAZZI202139, CORLI2022474, ELLIOTT1985257, KIM20151889].
Figure 1: Manifestation of the staircasing phenomenon widely observed in discretizations of the Perona-Malik equation, where the graphs of $u(\cdot,t)$ and $u_{x}(\cdot,t)$ are plotted at time $t=2$.
A natural approach to handling ill-posed problems is to introduce suitable regularization, often motivated by physical considerations, which leads to well-posed problems [Guidotti2015131]. Nevertheless, the qualitative behavior of solutions is highly sensitive to the type of regularization employed. In recent years, various regularization schemes have been investigated in different contexts (see [BELLETTINI2008892, Giacomelli2019374, GUIDOTTI20094731, Shao20221779]). The pseudo-parabolic regularization of equation (7),
$u_{t}=\left[\varPhi(u_{x})\right]_{x}+\varepsilon\left[\psi(u_{x})\right]_{xt},$ (9)
with $\varepsilon>0$ a small parameter and $\psi^{\prime}>0$, has been proposed to account for time-delay effects [Barenblatt1993341] and is applied to modeling the formation of layers with constant temperature or salinity in the ocean [Barenblatt19931414]. By the transformation $w=u_{x}$, equation (9) is formally related to
$w_{t}=\left[\varPhi(w)\right]_{xx}+\varepsilon\left[\psi(w)\right]_{xxt},$ (10)
which arises in population dynamics [Horstmann2004545]. Equation (10) serves as the pseudo-parabolic regularization of equation (8). If $\psi(u)=u$, this regularization is referred to as the Sobolev regularization [Victor1998457]. The well-posedness of the Neumann IBVP for (9) was established in [Barenblatt19931414], where $\psi$ is an increasing but bounded function. It was also shown that solutions may develop spatial discontinuities (jumps) within finite time, with amplitudes that increase over time. This phenomenon results not only from the shape of $\varPhi$ but also from the boundedness of $\psi$, whereas it does not occur if $\psi$ has power-type or logarithmic growth at infinity [Bertsch20131719, Bertsch2016051]. The existence of suitably defined nonnegative solutions to the Dirichlet IBVP for (10) was proved in the space of Radon measures (see [BERTSCH2015217, BERTSCH2016190, BERTSCH20172037, BERTSCH201846]), and results concerning uniqueness can be found in [Smarrazzo2012085]. As $\varepsilon\to 0^{+}$, the family of solutions to the regularized problem is expected to converge to a measure-valued solution of the original IBVP for (8); see [Smarrazzo2013551].
Both Dirichlet and Neumann IBVPs for the Cahn-Hilliard type regularization of equation (6),
$u_{t}=\operatorname{div}\vec{\varPhi}(\nabla u)-\varepsilon\Delta^{2}u,$ (11)
where $\vec{\varPhi}$ satisfies appropriate growth conditions, were investigated in [Slemrod1989001] (see also [Plotnikov1997550]). By the theory of Young measures (e.g., see [ball2005version, Giaquinta1998, Valadier1990]), it has been proved that a subsequence of the family of approximating solutions to the regularized problem (11) converges, in a suitable topology, to a Young measure solution of the ill-posed original IBVPs for (6). When the growth assumptions on $\vec{\varPhi}$ are relaxed, the Dirichlet IBVP for equation (6) was studied in [THANH20141403] through the limit $\varepsilon\rightarrow 0^{+}$ of solutions to the corresponding IBVP for Sobolev regularization
$u_{t}=\operatorname{div}\vec{\varPhi}(\nabla u)+\varepsilon\Delta u_{t},$ (12)
where the term $\Delta u_{t}$ can be interpreted as accounting for viscous relaxation effects [Novick1991331]. Specifically, the first step is to address equation (12) under the Dirichlet boundary condition for any $\varepsilon>0$, and then to examine the vanishing viscosity limit $\varepsilon\rightarrow 0^{+}$ of the family $\{u^{\varepsilon}\}_{\varepsilon}$ of solutions to the approximating problems, obtaining Young measure solutions for the original problem. The asymptotic behavior of such solutions for large times is also studied in [THANH20141403] using compactness arguments and $\omega$-limit set techniques. In view of the above results, a widely accepted strategy for handling forward-backward parabolic problems is to approximate them by a family of regularized problems, whose solutions are then used to extract limiting points, yielding suitably defined solutions of the original problem (see [Evans20041599, Oliver20031380, Smarrazzo20101046, Smarrazzo2013551]). It should be noted that, except in specific cases (see [Demoulini1996376, Mascia2009887]), the uniqueness of such solutions remains an open problem.
Figure 2: Microramping phenomenon observed for the regularized equation (1) with constant exponent $p$ and $\delta=0.001$.
Here, we propose a novel class of regularizations (1) for the Perona-Malik equation, interpretable as a $p(x)$-Laplacian regularization of forward-backward type, where the degree of regularization is adaptively modulated by spatial information. From the energy perspective, this regularization benefits from the variational structure
$\int_{\Omega}\left(\frac{1}{2}\log\left(1+\left|\nabla u\right|^{2}\right)+\frac{\delta}{p(x)}\left|\nabla u\right|^{p(x)}\right)dx,$ (13)
which provides a mild variational regularization of the convex-concave energy functional underlying the Perona-Malik flow ($\delta=0$), while any $\delta>0$ is sufficient to make (13) convex-concave-convex. Figures 13 illustrate the transition from rough to smooth staircasing. In contrast to (1) with a constant exponent, the solution to the proposed regularization for a suitable choice of $p$ shows better regularity; its gradients exhibit fewer oscillations, and the maximum gradients are confined to a smaller bounded interval. Meanwhile, equation (1) can be regarded as an independent anisotropic diffusion model in image processing, where the spatially varying exponent $p$ is designed according to practical requirements. In this sense, the new model serves as an interpolation between two classical models: the Perona-Malik model and that introduced in [CHEN20061383]. However, its application to image processing is not the primary focus of the present paper; interested readers are referred to [CHEN20061383, Guo2012958, Li20243374, SHAO2020103166] for further details on adaptive diffusion-based and variational models.
Figure 3: Gentler microramping phenomenon observed for the new regularized equation (1) with $\delta=0.001$.
To our knowledge, only a few papers address forward-backward parabolic problems with nonstandard growth conditions. Inspired by the idea in [Oliver20031380, THANH20141403], in this paper we study problem (1)–(3) using Sobolev regularization. Namely, we first investigate the existence and uniqueness of weak solutions to the homogeneous Neumann IBVP for the regularized equation
$u_{t}=\operatorname{div}\left(\left(\frac{1}{1+\left|\nabla u\right|^{2}}+\delta\left|\nabla u\right|^{p(x)-2}\right)\nabla u\right)+\varepsilon\Delta u_{t},$
where $\varepsilon\in(0,1)$. We then analyze the limit as $\varepsilon\rightarrow 0^{+}$ of the family $\{u^{\varepsilon}\}_{\varepsilon}$ of weak solutions to the above regularized problems and establish the existence of Young measure solutions to the original problem (1)–(3); see Definition 4. The uniqueness of such a solution is left as an open problem. We emphasize here that, although the treatment in [Yin2003659] (corresponding to equation (1) with constant exponent $1<p<2$; see also [Demoulini1996376]) involves Young measures, the spirit of that approach, which does not rely on a regularization scheme motivated by physical considerations, differs from the present one. Indeed, the approximation procedure in [Yin2003659] combines the explicit time discretization for the evolution equation with variational methods used to solve the stationary problem obtained at each time step. The difficulties arising from the nonconvexity of the potential are addressed via relaxation theory (see [Kinderlehrer1992001] for details), whereas the feasibility of our approach is based on the observation that the discretized counterpart of the damping term $\Delta u_{t}$ provides the convexity of the stationary problem even when the corresponding potential is nonconvex (see Lemma 10). However, the existence of such solutions for problem (1)–(3) cannot be established following the method in [Yin2003659], and, consequently, our work also serves as a theoretical complement to the studies in [GUIDOTTI20123226, Oliver20031380, Yin2003659]. Our main result is stated as follows:
Theorem 1 (Existence of Young measure solutions)
Let $f\in W^{1,p(\cdot)}(\Omega)$. Then problem (1)–(3) admits at least one Young measure solution.
The rest of this paper is organized as follows. Section 2 recalls some preliminaries on variable exponent spaces and Young measures, and introduces a family of regularized problems approximating problem (1)–(3), together with the definition of Young measure solutions. Section 3 establishes the existence and uniqueness of Young measure solutions to the regularized problem. Finally, Section 4 presents the proof of Theorem 1.
2 Preliminaries
In this section, we present some definitions and results concerning Lebesgue and Sobolev spaces with variable exponents, as well as Young measures. We also define a Young measure solution to problem (1)–(3) and establish a family of corresponding regularized problems.
2.1 Variable exponent spaces
We first introduce the notations for variable exponent spaces. Let $\Omega$ be an open subset of $\mathbb{R}^{N}$. We denote by $\mathcal{P}(\Omega)$ the set of all Lebesgue measurable functions $p:\Omega\rightarrow[1,\infty]$. Functions $p\in\mathcal{P}(\Omega)$ are called variable exponents on $\Omega$. If $p^{+}<\infty$, we say that $p$ is a bounded variable exponent. For any $p\in\mathcal{P}(\Omega)$, we define $p^{\prime}\in\mathcal{P}(\Omega)$ by $\frac{1}{p(y)}+\frac{1}{p^{\prime}(y)}=1$, where $\frac{1}{\infty}:=0$. The function $p^{\prime}$ is called the dual variable exponent of $p$.
Let $p\in\mathcal{P}(\Omega)$. We define the semimodular
$\varrho_{L^{p(\cdot)}(\Omega)}(u):=\int_{\Omega}\left|u(x)\right|^{p(x)}dx.$
The variable exponent Lebesgue space $L^{p(\cdot)}(\Omega)$ is given by
$L^{p(\cdot)}(\Omega)=\left\{u:\Omega\rightarrow\mathbb{R}\text{ measurable},\varrho_{L^{p(\cdot)}(\Omega)}(\lambda u)<\infty\text{ for some }\lambda>0\right\},$
equipped with the Luxemburg norm
$\left\|u\right\|_{L^{p(\cdot)}(\Omega)}=\inf\left\{\lambda>0:\varrho_{L^{p(\cdot)}(\Omega)}\left(\frac{u}{\lambda}\right)\leq 1\right\}.$
The function $u\in L^{p(\cdot)}(\Omega)$ belongs to the space $W^{k,p(\cdot)}(\Omega)$, where $k\in\mathbb{N}$ and $p\in\mathcal{P}(\Omega)$, if its weak partial derivatives $\partial_{\alpha}u$ with $\left|\alpha\right|\leq k$ exist and lie in $L^{p(\cdot)}(\Omega)$. We define the Sobolev semimodular on $W^{k,p(\cdot)}(\Omega)$ by
$\varrho_{W^{k,p(\cdot)}(\Omega)}(u):=\sum_{0\leq\left|\alpha\right|\leq k}\varrho_{L^{p(\cdot)}(\Omega)}(\partial_{\alpha}u),$
which induces the norm
$\left\|u\right\|_{W^{k,p(\cdot)}(\Omega)}=\inf\left\{\lambda>0:\varrho_{W^{k,p(\cdot)}(\Omega)}\left(\frac{u}{\lambda}\right)\leq 1\right\}.$
Furthermore, the variable exponent Sobolev space $W^{k,p(\cdot)}(\Omega)$ admits the equivalent norm
$\sum_{m=0}^{k}\left\|\nabla^{m}u\right\|_{L^{p(\cdot)}(\Omega)}=\sum_{m=0}^{k}\left\|\left|\nabla^{m}u\right|\right\|_{L^{p(\cdot)}(\Omega)}.$
We recall the log-Hölder continuity condition, which is widely used in the study of variable exponent spaces.
Definition 1
A function $\beta:\Omega\rightarrow\mathbb{R}$ is called locally $\log$-Hölder continuous on $\Omega$ if there exists $a_{1}>0$ such that
$\left|\beta(x)-\beta(y)\right|\leq\frac{a_{1}}{\log\left(e+1/\left|x-y\right|\right)}$
for all $x,y\in\Omega$. A function $\beta$ is said to satisfy the log-Hölder decay condition if there exist $\beta_{\infty}\in\mathbb{R}$ and a constant $a_{2}>0$ such that
$\left|\beta(x)-\beta_{\infty}\right|\leq\frac{a_{2}}{\log\left(e+\left|x\right|\right)}$
for all $x\in\Omega$. We say that $\beta$ is globally $\log$-Hölder continuous in $\Omega$ if it is locally log-Hölder continuous and satisfies the log-Hölder decay condition. The constants $a_{1}$ and $a_{2}$ are referred to as the local log-Hölder constant and the log-Hölder decay constant, respectively. The maximum $\max\{a_{1},a_{2}\}$ is called the log-Hölder constant of $\beta$.
We introduce the following class of variable exponents
[…]
The $\log$-Hölder constant of $\frac{1}{p}$ is denoted by $c_{\log}(p)$ or $c_{\log}$. Moreover, if $p\in\mathcal{P}(\Omega)$ with $p^{+}<\infty$, then $p\in\mathcal{P}^{\log}(\Omega)$ if and only if $p$ is globally $\log$-Hölder continuous.
Definition 2
A bounded domain $\Omega\subset\mathbb{R}^{N}$ is called an $\alpha$-John domain, with $\alpha>0$, if there exists $x_{0}\in\Omega$ such that each point in $\Omega$ can be joined to $x_{0}$ by a rectifiable path $\gamma$, parametrized by arc-length, satisfying
$B\left(\gamma(t),\frac{1}{\alpha}t\right)\subset\Omega$
for all $t\in[0,\ell(\gamma)]$, where $\ell(\gamma)$ is the length of $\gamma$. The ball $B\left(x_{0},\frac{1}{2\alpha}\operatorname{diam}(\Omega)\right)$ is called the John ball. If the value of $\alpha$ is irrelevant, we refer to $\Omega$ as a John domain.
There are many examples of John domains. Every ball is clearly a 1-John domain, and every bounded domain satisfying the uniform interior cone condition is a John domain. In particular, any bounded Lipschitz domain is a John domain, where a domain with $C^{0,1}$-boundary is referred to as a Lipschitz domain; see also [Diening2011, Martio1979383].
For later use, we state the following important properties of the spaces $L^{p(\cdot)}(\Omega)$ and $W^{1,p(\cdot)}(\Omega)$. Detailed discussions can be found in [Diening2011, Kovavcik1991592, Mihuailescu20072929].
Lemma 1
[Diening2011] Let $p\in\mathcal{P}(\Omega)$ with $p^{-}<\infty$. If $\varrho_{L^{p(\cdot)}(\Omega)}(u)>0$ or $p^{+}<\infty$, then
$\min\left\{\left\|u\right\|_{L^{p(\cdot)}(\Omega)}^{p^{-}},\left\|u\right\|_{L^{p(\cdot)}(\Omega)}^{p^{+}}\right\}\leq\varrho_{L^{p(\cdot)}(\Omega)}(u)\leq\max\left\{\left\|u\right\|_{L^{p(\cdot)}(\Omega)}^{p^{-}},\left\|u\right\|_{L^{p(\cdot)}(\Omega)}^{p^{+}}\right\}.$
Lemma 2
[Diening2011] Let $p\in\mathcal{P}(\Omega)$. Then $L^{p(\cdot)}(\Omega)$ and $W^{k,p(\cdot)}(\Omega)$ are Banach spaces, separable if $p$ is bounded, and reflexive and uniformly convex whenever $1<p^{-}\leq p^{+}<\infty$.
Lemma 3 (Hölder’s inequality)
[Diening2011] Let $p,q,s\in\mathcal{P}(\Omega)$ be such that $\frac{1}{s(y)}=\frac{1}{p(y)}+\frac{1}{q(y)}$ for a.e. $y\in\Omega$. Then, for all $u\in L^{p(\cdot)}(\Omega)$ and $v\in L^{q(\cdot)}(\Omega)$,
$\left\|uv\right\|_{L^{s(\cdot)}(\Omega)}\leq 2\left\|u\right\|_{L^{p(\cdot)}(\Omega)}\left\|v\right\|_{L^{q(\cdot)}(\Omega)},$
where in the case $s=p=q=\infty$, we adopt the convention $\frac{s}{p}=\frac{s}{q}=1$.
Lemma 4
[Diening2011] Let $\Omega$ be a Lipschitz domain, and let $p\in\mathcal{P}^{\log}(\Omega)$ satisfy $1\leq p^{-}\leq p^{+}<\infty$. Then $C^{\infty}(\bar{\Omega})$ is dense in $W^{k,p(\cdot)}(\Omega)$, $k\in\mathbb{N}$.
Lemma 5 (Poincaré inequality)
[Diening2011] Let $p\in\mathcal{P}^{\log}(\Omega)$. If $\Omega$ is a bounded $\alpha$-John domain, then
$\left\|u-\frac{1}{\left|\Omega\right|}\int_{\Omega}udx\right\|_{L^{p(\cdot)}(\Omega)}\leq c\operatorname{diam}(\Omega)\left\|\nabla u\right\|_{L^{p(\cdot)}(\Omega)},$
for $u\in W^{1,p(\cdot)}(\Omega)$. The constant $c$ depends only on the dimension $N$, $\alpha$ and $c_{\log}(p)$.
Lemma 6
Let $p\in\mathcal{P}(\Omega)$ with $1<p^{-}\leq p^{+}<\infty$. Suppose $\{f^{m}\}_{m=1}^{\infty}\subset L^{\infty}(0,T;L^{p(\cdot)}(\Omega))$ and $f^{m}\rightharpoonup 0$ weakly $\ast$ in $L^{\infty}(0,T;L^{p(\cdot)}(\Omega))$ as $m\rightarrow\infty$. Let $\varPsi$ be a compact subset of $L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))$. Then
$\sup_{\xi\in\varPsi}\left|\iint_{Q_{T}}f^{m}\xi dxdt\right|\rightarrow 0,\quad\text{as }m\rightarrow\infty.$
Proof 1
Here we may assume that $\left\|f^{m}\right\|_{L^{\infty}(0,T;L^{p(\cdot)}(\Omega))}\leq C$ for all $m\in\mathbb{N}$. Since $\varPsi\subset L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))$ is compact, it is totally bounded. Given any $r>0$, there exists a finite subset $\left\{\xi_{1},\xi_{2},\cdots,\xi_{n}\right\}\subset\varPsi$ such that, for every $\eta\in\varPsi$, one can find $i\in\{1,2,\cdots,n\}$ satisfying
$\left\|\eta-\xi_{i}\right\|_{L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))}<r.$
By the weak $\ast$ convergence of $\{f^{m}\}_{m=1}^{\infty}$ (see [Zeidler1990]), there exists $N_{i}\in\mathbb{N}$ such that, for all $m>N_{i}$, it holds that
$\left|\iint_{Q_{T}}f^{m}\xi_{i}dxdt\right|<r,\quad i=1,2,\cdots,n.$
Fix $N=\max_{1\leq i\leq n}N_{i}$. For any $\xi\in\varPsi$, there exists $j\in\{1,2,\cdots,n\}$ such that $\left\|\xi-\xi_{j}\right\|_{L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))}<r$. Consequently, for all $m>N$, it follows from Lemma 3 that
$\left|\iint_{Q_{T}}f^{m}\xi dxdt\right|\leq$ $\left|\iint_{Q_{T}}f^{m}\xi_{j}dxdt\right|+\left|\int_{0}^{T}\int_{\Omega}f^{m}\left(\xi-\xi_{j}\right)dxdt\right|$
$\leq$ $r+2\int_{0}^{T}\left\|f^{m}(t)\right\|_{L^{p(\cdot)}(\Omega)}\left\|\xi(t)-\xi_{j}(t)\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}dt$
$\leq$ $r+2\left\|f^{m}\right\|_{L^{\infty}(0,T;L^{p(\cdot)}(\Omega))}\left\|\xi-\xi_{j}\right\|_{L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))}$
$\leq$ $\left(1+2C\right)r.$
This completes the proof.
Adapting the argument of Theorem 3.20 in [Dacorogna2008], the following sufficient condition for weak lower semicontinuity on $W^{1,p(\cdot)}(\Omega)$ can be derived. The proof is omitted for brevity.
Lemma 7
Assume $p\in\mathcal{P}(\Omega)$ with $1\leq p^{-}\leq p^{+}<\infty$, and let $f:\Omega\times\mathbb{R}^{N}\rightarrow$ $\mathbb{R}\cup\{+\infty\}$ be a Carathéodory function satisfying
$f(x,\xi)\geq a(x)\cdot\xi+b(x)$
for a.e. $x\in\Omega$ and for all $\xi\in\mathbb{R}^{N}$, where $a\in L^{p^{\prime}(\cdot)}(\Omega;\mathbb{R}^{N})$ and $b\in L^{1}(\Omega)$. Let
$J(\xi):=\int_{\Omega}f(x,\xi(x))dx.$
Assume that $\xi\rightarrow f(x,\xi)$ is convex and that
$\xi_{\nu}\rightharpoonup\xi\text{ weakly in }L^{p(\cdot)}(\Omega).$
Then
$\liminf_{\nu\rightarrow\infty}J\left(\xi_{\nu}\right)\geq J(\xi).$
2.2 Young measures
For the readers’ convenience, we briefly recall some fundamental notions concerning Young measures. Denote by $C_{0}(\mathbb{R}^{N})$ the closure of continuous functions in $\mathbb{R}^{N}$ with compact support. The dual of $C_{0}(\mathbb{R}^{N})$ can be identified with the space $\mathcal{M}(\mathbb{R}^{N})$ of signed Radon measures with finite mass, via the pairing
$\langle\mu,f\rangle=\int_{\mathbb{R}^{N}}f(\xi)d\mu(\xi),\quad f\in C_{0}(\mathbb{R}^{N}),~\mu\in\mathcal{M}(\mathbb{R}^{N}).$
A map $\nu:\Omega\rightarrow\mathcal{M}(\mathbb{R}^{N})$ is called weakly $\ast$ measurable if the functions $x\mapsto\langle\nu_{x},f\rangle$ are measurable for all $f\in C_{0}(\mathbb{R}^{N})$, where $\nu_{x}=\nu(x)$. We state the following fundamental theorem on Young measures, as given by J. M. Ball.
Lemma 8
[ball2005version] Let $\Omega\subset\mathbb{R}^{N}$ be Lebesgue measurable, let $K\subset\mathbb{R}^{N}$ be closed, and let $u_{j}:\Omega\rightarrow\mathbb{R}^{N},j\in\mathbb{N}$, be a sequence of Lebesgue measurable functions satisfying $u_{j}\rightarrow K$ in measure as $j\rightarrow\infty$, i.e., given any open neighbourhood $U$ of $K$ in $\mathbb{R}^{N}$
$\lim_{j\rightarrow\infty}\left|\left\{x\in\Omega:u_{j}(x)\notin U\right\}\right|=0.$
Then there exists a subsequence $\{u_{j_{k}}\}_{k=1}^{\infty}$ of $\{u_{j}\}_{j=1}^{\infty}$ and a weakly $\ast$ measurable map $\nu:\Omega\rightarrow\mathcal{M}(\mathbb{R}^{N})$ such that
  • (i)
    $\nu(x)\geq 0$ and $\|\nu(x)\|_{\mathcal{M}\left(\mathbb{R}^{N}\right)}=\int_{\mathbb{R}^{N}}d\nu_{x}\leq 1$ for a.e. $x\in\Omega$;
  • (ii)
    $\operatorname{spt}\nu_{x}\subset K$ for a.e. $x\in\Omega$;
  • (iii)
    For all $\psi\in C_{0}(\mathbb{R}^{N})$, $\psi(u_{j_{k}})\rightharpoonup\langle\nu,\psi\rangle$ weakly $\ast$ in $L^{\infty}(\Omega)$.
Suppose further that $\{u_{j_{k}}\}_{k=1}^{\infty}$ satisfies the boundedness condition
$\forall R>0:~\lim_{L\rightarrow\infty}\sup_{k\in\mathbb{N}}\left|\{x\in\Omega\cap B_{R}:|u_{j_{k}}(x)|\geq L\}\right|=0,$ (14)
where $B_{R}=B_{R}(0)$. Then
$\left\|\nu_{x}\right\|_{\mathcal{M}(\mathbb{R}^{N})}=1\quad\text{ for a.e. }x\in\Omega,$ (15)
and the following condition holds
[…]
Definition 3
The map $\nu:\Omega\rightarrow\mathcal{M}(\mathbb{R}^{N})$ in Lemma 8 is called the Young measure in $\mathbb{R}^{N}$ generated by the sequence $\{u_{j_{k}}\}_{k=1}^{\infty}$.
Hungerbühler subsequently established the following consequence, while refined versions of Lemma 8 can be found in [Kinderlehrer1994059].
Lemma 9
[Hungerbuhler1997] Let $\Omega$, $u_{j}$, and $\nu_{x}$ be as in Lemma 8. Then (14), (15) and (16) are equivalent.
Remark 1
In [ball2005version], it is also shown that, under hypothesis (14), for any measurable set $D\subset\Omega$,
[…]
for every Carathéodory function $f:D\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ such that $\left\{f(\cdot,u_{j_{k}})\right\}_{k=1}^{\infty}$ is sequentially weakly relatively compact in $L^{1}(D)$. This statement is thus also equivalent to (14), (15), and (16).
2.3 Definition of Young measure solutions
We define a Young measure solution to problem (1)–(3). For clarity, let $\rho\in W^{1,p(\cdot)}(\Omega)$ be fixed but arbitrary, and define
$W_{\rho}^{1,p(\cdot)}(\Omega)=\left\{u\in W^{1,p(\cdot)}(\Omega):\int_{\Omega}udx=\int_{\Omega}\rho dx\right\}$
as the affine subspace of $W^{1,p(\cdot)}(\Omega)$. Recall that the variable exponent $p$ satisfies condition (4), and denote by
$\vec{q}(x,\xi)=\frac{\xi}{1+\left|\xi\right|^{2}}+\delta\left|\xi\right|^{p(x)-2}\xi,\quad x\in\Omega,~\xi\in\mathbb{R}^{N},$
the heat flux in equation (1), where $\vec{q}(x,\xi)=\nabla_{2}\varphi(x,\xi)$ and $\nabla_{2}$ denotes the derivative with respect to the second variable. The potential $\varphi:\Omega\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ is a Carathéodory function, nonconvex with respect to $\xi$ for sufficiently small $\delta$. It is noted that $\varphi(x,\cdot)\in C^{2}(\mathbb{R}^{N})$ for all $x\in\Omega$, and that $\vec{q}$ and $\varphi$ satisfy the following growth conditions:
$\left|\vec{q}(x,\xi)\right|\leq\lambda_{2}\left|\xi\right|^{p(x)-1}+1,\quad\max\left\{\lambda_{1}\left|\xi\right|^{p(x)}-1,0\right\}\leq\varphi(x,\xi)\leq\lambda_{2}\left|\xi\right|^{p(x)}+1,$ (17)
for all $x\in\Omega$, $\xi\in\mathbb{R}^{N}$, where $0<\lambda_{1}\leq\lambda_{2}$ are constants. Furthermore, there exists a constant $K=K(N)>0$ such that
$\left(\vec{q}(x,\xi_{1})-\vec{q}(x,\xi_{2})\right)\cdot\left(\xi_{1}-\xi_{2}\right)\geq-K\left|\xi_{1}-\xi_{2}\right|^{2},$ (18)
for all $x\in\Omega$, $\xi_{1},\xi_{2}\in\mathbb{R}^{N}$.
Definition 4
A pair $(u,\nu)$ is called a Young measure (or measure-valued) solution of problem (1)–(3) if $u\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;L^{2}(\Omega))$ and $\nu=(\nu_{x,t})_{(x,t)\in Q_{T}}$ is a parametrized family of probability measures on $\mathbb{R}^{N}$, such that $u(0)=f$ a.e. in $\Omega$, and
$\nabla u(x,t)=\langle\nu_{x,t},\mathrm{id}\rangle=\int_{\mathbb{R}^{N}}\xi d\nu_{x,t}(\xi)\quad\text{a.e. in }Q_{T},$ (19)
$\iint_{Q_{T}}\left(u_{t}\zeta+\langle\nu_{x,t},\vec{q}(x,\cdot)\rangle\cdot\nabla\zeta\right)dxdt=0,$ (20)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$, where $\mathrm{id}$ denotes the unit mapping in $\mathbb{R}^{N}$.
Our goal is to construct a Young measure solution to problem (1)–(3). Following the approach in [Oliver20031380, THANH20141403], we first consider the homogeneous Neumann IBVP for the Sobolev regularization of equation (1):
$\displaystyle u^{\varepsilon}_{t}=\operatorname{div}\vec{q}(x,\nabla u^{\varepsilon})+\varepsilon\Delta u^{\varepsilon}_{t},$ $(x,t)\in Q_{T}$, (21)
$\displaystyle\left(\vec{q}\left(x,\nabla u^{\varepsilon}\right)+\varepsilon\nabla u_{t}^{\varepsilon}\right)\cdot\vec{n}=0,$ $(x,t)\in\partial\Omega\times(0,T)$, (22)
$\displaystyle u^{\varepsilon}(x,0)=f,$ $x\in\Omega$, (23)
where $0<\varepsilon<1$. We refer to the problem (21)–(23) as the regularized problem corresponding to the original problem (1)–(3).
3 Weak solutions of the regularized problem (21)–(23)
In this section, we study the well-posedness of problem (21)–(23) and establish the following theorem by employing Rothe’s method and variational principles.
Theorem 2
Let $f\in W^{1,p(\cdot)}(\Omega)$. Then the regularized problem (21)–(23) admits a unique weak solution $u^{\varepsilon}\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$, satisfying $u^{\varepsilon}(0)=f$ a.e. in $\Omega$, and
$\iint_{Q_{T}}\left(\frac{\partial u^{\varepsilon}}{\partial t}\zeta+\vec{q}(x,\nabla u^{\varepsilon})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u^{\varepsilon}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (24)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Furthermore, for a.e. $t\in[0,T]$, the following inequality holds:
$\int_{\Omega}\varphi(x,\nabla u^{\varepsilon}(t))dx+\left\|\frac{\partial u^{\varepsilon}}{\partial\tau}\right\|^{2}_{L^{2}(Q_{t})}+\left\|\nabla\sqrt{\varepsilon}\frac{\partial u^{\varepsilon}}{\partial\tau}\right\|^{2}_{L^{2}(Q_{t})}\leq\int_{\Omega}\varphi(x,\nabla f)dx,$ (25)
where $Q_{t}=\Omega\times(0,t)$.
Remark 2
By substituting the constant test function $\zeta=1_{\Omega}$ into the weak formulation (25), we obtain
$\frac{d}{dt}\int_{\Omega}u^{\varepsilon}dx=0,$
which implies that the mean of any solution is conserved, i.e.,
$\int_{\Omega}u^{\varepsilon}(x,t)dx=\int_{\Omega}u^{\varepsilon}(x,0)dx=\int_{\Omega}f(x)dx.$
Therefore, it is natural to require that the solution to problem (21)–(23) satisfies $u^{\varepsilon}(\cdot,t)\in W_{f}^{1,p(\cdot)}(\Omega)$ for a.e. $t\in[0,T]$.
For simplicity, we omit the superscript $\varepsilon$ in the notation $u^{\varepsilon}$ and write $u$ instead. Let $m$ be a positive integer. For $u\in W^{1,p(\cdot)}(\Omega)$, we define
$J^{m,j}(u)=\int_{\Omega}\left(\varphi(x,\nabla u)+\frac{m}{2T}\left|u-u_{m}^{j-1}\right|^{2}+\varepsilon\frac{m}{2T}\left|\nabla u-\nabla u_{m}^{j-1}\right|^{2}\right)dx,$
where $u_{m}^{0}=f$, $u_{m}^{j-1}\in W_{f}^{1,p(\cdot)}(\Omega)$, $j=1,2,\cdots,m$. The energy functional above is lower bounded due to the growth condition (17). It is readily seen that for every $v\in W^{1,p(\cdot)}(\Omega)$,
$J^{m,1}\left(v-\frac{1}{\left|\Omega\right|}\int_{\Omega}vdx+\frac{1}{\left|\Omega\right|}\int_{\Omega}fdx\right)\leq J^{m,1}(v).$
Without loss of generality, minimizers and minimizing sequences can be assumed to lie in $W_{f}^{1,p(\cdot)}(\Omega)$. First, we prove the following lemma.
Lemma 10
If $m$ is sufficiently large and $u_{m}^{j-1}\in W_{f}^{1,p(\cdot)}(\Omega)$, then $J^{m,j}$ admits a minimizer $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ for $j=1,2,\cdots,m$.
Proof 2
It is worth mentioning that the term in $J^{m,j}$ involving the discretized counterpart of $\nabla u_{t}$ ensures the convexity of the integrand with respect to $\nabla u$, although $\varphi$ itself is not convex. Introducing the notation
$\varphi_{\varepsilon}(x,\xi)=\varphi(x,\xi)+\varepsilon\frac{m}{2T}\left|\xi\right|^{2}$
and rewriting
$J^{m,j}(u)=\int_{\Omega}\left(\varphi_{\varepsilon}(x,\nabla u)+\frac{m}{2T}\left|u-u_{m}^{j-1}\right|^{2}-\varepsilon\frac{m}{T}\nabla u\cdot\nabla u_{m}^{j-1}+\varepsilon\frac{m}{2T}\left|\nabla u_{m}^{j-1}\right|^{2}\right)dx.$
It is clear that $u\mapsto J^{m,j}(u)-\int_{\Omega}\varphi_{\varepsilon}(x,\nabla u)dx$ is weakly lower semicontinuous on $W^{1,p(\cdot)}(\Omega)$. Additionally, by (18),
$\left(\nabla_{2}\varphi_{\varepsilon}(x,\xi_{1})-\nabla_{2}\varphi_{\varepsilon}(x,\xi_{2})\right)\cdot\left(\xi_{1}-\xi_{2}\right)\geq\left(-K+\varepsilon\frac{m}{T}\right)\left|\xi_{1}-\xi_{2}\right|^{2},\quad x\in\Omega,~\xi_{1},\xi_{2}\in\mathbb{R}^{N},$ (26)
which shows that the mapping $\xi\mapsto\varphi_{\varepsilon}(x,\xi)$ is convex for every $x\in\Omega$ whenever $m\geq\frac{KT}{\varepsilon}$. Lemma 7 then yields the weak lower semicontinuity of $u\mapsto\int_{\Omega}\varphi_{\varepsilon}(x,\nabla u)dx$ on $W^{1,p(\cdot)}(\Omega)$. Let $\{u_{m}^{j,k}\}_{k=1}^{\infty}\subset W_{f}^{1,p(\cdot)}(\Omega)$ be a minimizing sequence of $J^{m,j}$. Since $J^{m,j}$ is coercive on $W_{f}^{1,p(\cdot)}(\Omega)$ by Lemma 5, we may assume that $\{u_{m}^{j,k}\}_{k=1}^{\infty}$ is bounded in $W^{1,p(\cdot)}(\Omega)$. Applying Lemma 2, there exists $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ and a subsequence of $\{u_{m}^{j,k}\}_{k=1}^{\infty}$, still denoted by itself, such that $u_{m}^{j,k}\rightharpoonup u_{m}^{j}$ weakly in $W^{1,p(\cdot)}(\Omega)$. Then, we have
$\inf_{u\in W^{1,p(\cdot)}(\Omega)}J^{m,j}(u)\leq J^{m,j}(u_{m}^{j})\leq\liminf_{k\rightarrow\infty}J^{m,j}(u_{m}^{j,k})=\inf_{u\in W^{1,p(\cdot)}(\Omega)}J^{m,j}(u).$
Consequently, $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ is a (not necessarily unique) minimizer of $J^{m,j}$. This finishes the proof.
Note that, for $j=1,2,\cdots,m$, the minimizers $u_{m}^{j}$ satisfy the following Euler-Lagrange equations for every $\phi\in C^{\infty}(\bar{\Omega})$,
$\int_{\Omega}\left(\frac{m}{T}\left(u_{m}^{j}-u_{m}^{j-1}\right)\phi+\vec{q}(x,\nabla u_{m}^{j})\cdot\nabla\phi+\varepsilon\frac{m}{T}\left(\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right)\cdot\nabla\phi\right)dx=0.$ (27)
From (17) and the facts that $\nabla u_{m}^{j}\in L^{p(\cdot)}(\Omega)$ and $\vec{q}(x,\nabla u_{m}^{j})\in L^{p^{\prime}(\cdot)}(\Omega)$, together with Lemma 4, it follows that $\phi\in W^{1,p(\cdot)}(\Omega)$ is admissible as a test function in the the Euler-Lagrange system (27) for $u_{m}^{j}$. We now derive an a priori estimate for the discrete energy.
Lemma 11 (Discrete energy inequality)
Given $\gamma\in(0,\varepsilon)$, there exists a constant $m_{0}(\gamma)>0$, such that for all $m\geq m_{0}$, the minimizers $u_{m}^{j}$ of the functionals $J^{m,j}$ satisfy the estimate
$\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx+\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx+\sum_{i=1}^{j}\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx\leq\int_{\Omega}\varphi(x,\nabla f)dx,$ (28)
where $j=1,2,\cdots,m$.
Proof 3
Similar to the treatment in (26), it follows from (18) that, for $m\geq\frac{KT}{2\gamma}$, the mapping $\xi\mapsto\varphi(x,\xi)+\gamma\frac{m}{T}|\xi-\nabla u_{m}^{j-1}|^{2}$, $j=1,2,\cdots,m$, is convex; i.e., for all $x\in\Omega$ and $\xi_{1},\xi_{2}\in\mathbb{R}^{N}$, we have
$\varphi(x,\xi_{1})+$ $\gamma\frac{m}{T}\left|\xi_{1}-\nabla u_{m}^{j-1}(x)\right|^{2}-\left(\varphi(x,\xi_{2})+\gamma\frac{m}{T}\left|\xi_{2}-\nabla u_{m}^{j-1}(x)\right|^{2}\right)$
$\leq\left(\vec{q}(x,\xi_{1})+2\gamma\frac{m}{T}\left(\xi_{1}-\nabla u_{m}^{j-1}(x)\right)\right)\cdot\left(\xi_{1}-\xi_{2}\right).$ (29)
Set $m_{0}=\max\left\{\frac{KT}{2\gamma},\frac{KT}{\varepsilon}\right\}$. Then, by Lemma 10, $J^{m,j}$ admits a minimizer $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$, satisfying (27) for every $m\geq m_{0}$ and $j=1,2,\cdots,m$. From (29), for $i=1,2,\cdots,j$, we derive
$\int_{\Omega}\varphi(x,\nabla u_{m}^{i})dx-\int_{\Omega}\varphi(x,\nabla u_{m}^{i-1})dx$
$=$ $\int_{\Omega}\bigg[\varphi(x,\nabla u_{m}^{i})+\gamma\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}-\left(\varphi(x,\nabla u_{m}^{i-1})+\gamma\frac{m}{T}\left|\nabla u_{m}^{i-1}-\nabla u_{m}^{i-1}\right|^{2}\right)$
$-\gamma\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}\bigg]dx$
$\leq$ $\int_{\Omega}\bigg[\left(\vec{q}(x,\nabla u_{m}^{i})+2\gamma\frac{m}{T}\left(\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right)\right)\cdot\left(\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right)$
$-\gamma\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}\bigg]dx$
$=$ $\int_{\Omega}\bigg[\left(\vec{q}(x,\nabla u_{m}^{i})+\varepsilon\frac{m}{T}\left(\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right)\right)\cdot\left(\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right)$
$-(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}\bigg]dx.$ (30)
Before proceeding with the estimate, we consider (27) with $\phi=u_{m}^{i}-u_{m}^{i-1}$, which yields
$\int_{\Omega}\left(\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}+\left(\vec{q}(x,\nabla u_{m}^{i})+\varepsilon\frac{m}{T}\left(\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right)\right)\cdot\left(\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right)\right)dx=0.$
We insert the above equality into (30) to obtain
$\int_{\Omega}\varphi(x,\nabla u_{m}^{i})dx-\int_{\Omega}\varphi(x,\nabla u_{m}^{i-1})dx\leq-\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx-\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$
Summing the above inequality from $i=1$ to $j$ leads to
$\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx-\int_{\Omega}\varphi(x,\nabla f)dx\leq-\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx-\sum_{i=1}^{j}\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$
This completes the proof.
Corollary 1
For every $m\geq m_{0}(\gamma)$, the minimizers of the functionals $J^{m,j}$ further satisfy the following estimate
$\sup_{1\leq j\leq m}\left\|u_{m}^{j}\right\|_{W^{1,p(\cdot)}(\Omega)}+\sum_{j=1}^{m}\frac{m}{T}\left\|u_{m}^{j}-u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}+\sum_{j=1}^{m}\left(\varepsilon-\gamma\right)\frac{m}{T}\left\|\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}\leq M_{1},$ (31)
where $M_{1}>0$ depends only on $N$, $\Omega$, $\lambda_{1}$, $\lambda_{2}$, $p^{+}$, $c_{\log}(p)$, $f$, and $\alpha$, the John constant of $\Omega$.
The assertion follows directly from the growth condition (17), Lemma 5 and 11, and the assumption that $f\in W^{1,p(\cdot)}(\Omega)$.
Now we construct approximate solutions to the regularized problem (21)–(23). Let $m\geq m_{0}$. For $j=1,2,\cdots,m$, let $\chi_{m}^{j}(t)$ denote the characteristic function of the interval $(j-1)\frac{T}{m},j\frac{T}{m})$, and define $\lambda_{m}^{j}(t)$ as
$\lambda_{m}^{j}(t)=\left(\frac{m}{T}t-(j-1)\right)\chi_{m}^{j}(t),\quad 0\leq t\leq T.$
Define the following functions on $Q_{T}$
$u_{m}(x,t)=\sum_{j=1}^{m}\chi_{m}^{j}(t)\left\{u_{m}^{j-1}(x)+\lambda_{m}^{j}(t)\left(u_{m}^{j}(x)-u_{m}^{j-1}(x)\right)\right\},$
$\bar{u}_{m}(x,t)=\sum_{j=1}^{m}\chi_{m}^{j}(t)u_{m}^{j}(x).$
Then the Euler-Lagrange equation (27) can be equivalently rewritten as
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\zeta+\vec{q}(x,\nabla\bar{u}_{m})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (32)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. To establish the existence of weak solutions to problem (21)–(23), we use Corollary 1 to obtain uniform estimates for the approximate solutions.
Lemma 12
There exists a positive constant $M_{2}$, depending only on $M_{1}$, $\Omega$, $T$, $\lambda_{2}$, and $p$, such that the following inequalities hold:
$\sup_{0\leq t\leq T}\left\|u_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}\leq M_{2},~\left\|\frac{\partial u_{m}}{\partial t}\right\|_{L^{2}(Q_{T})}\leq M_{2},~\left\|\nabla\sqrt{\varepsilon-\gamma}\frac{\partial u_{m}}{\partial t}\right\|_{L^{2}(Q_{T})}\leq M_{2},$ (33)
$\sup_{0\leq t\leq T}\left\|\bar{u}_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}\leq M_{2},~\sup_{0\leq t\leq T}\left\|\vec{q}(x,\nabla\bar{u}_{m})\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}\leq M_{2},$ (34)
$\left\|u_{m}-\bar{u}_{m}\right\|_{L^{2}(Q_{T})}\leq\frac{1}{m}M_{2}.$ (35)
Proof 4
For any $t\in[0,T)$, there exists a positive integer $j=1,2,\cdots,m$ such that $(j-1)\frac{T}{m}\leq t<j\frac{T}{m}$. Given that $\lambda_{m}^{j}(t)\in[0,1)$, it follows from (31) that
$\left\|u_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}\leq\lambda_{m}^{j}(t)\left\|u_{m}^{j}\right\|_{W^{1,p(\cdot)}(\Omega)}+\left(1-\lambda_{m}^{j}(t)\right)\left\|u_{m}^{j-1}\right\|_{W^{1,p(\cdot)}(\Omega)}\leq M_{1},$
and, analogously, we obtain the corresponding estimate for $\bar{u}_{m}$. Furthermore, we derive from (17) and Lemma 1 that
$\varrho_{L^{p^{\prime}(\cdot)}(\Omega)}\left(\vec{q}(x,\nabla\bar{u}_{m}(t))\right)=$ $\int_{\Omega}\left|\vec{q}(x,\nabla u_{m}^{j})\right|^{\frac{p(x)}{p(x)-1}}dx$
$\leq$ $\int_{\Omega}\left(\lambda_{2}\left|\nabla u_{m}^{j}\right|^{p(x)-1}+1\right)^{\frac{p(x)}{p(x)-1}}dx$
$\leq$ $\int_{\Omega}2^{\frac{1}{p(x)-1}}\left(\lambda_{2}^{\frac{p(x)}{p(x)-1}}\left|\nabla u_{m}^{j}\right|^{p(x)}+1\right)dx$
$\leq$ $2^{\frac{1}{p^{-}-1}}(\lambda_{2}+1)^{\frac{p^{-}}{p^{-}-1}}\int_{\Omega}\left|\nabla u_{m}^{j}\right|^{p(x)}dx+2^{\frac{1}{p^{-}-1}}\left|\Omega\right|$
$\leq$ $2^{\frac{1}{p^{-}-1}}\left(\lambda_{2}+1\right)^{\frac{p^{-}}{p^{-}-1}}\max\left\{\left\|\nabla u_{m}^{j}\right\|_{L^{p(\cdot)}(\Omega)}^{p^{-}},\left\|\nabla u_{m}^{j}\right\|_{L^{p(\cdot)}(\Omega)}^{p^{+}}\right\}+2^{\frac{1}{p^{-}-1}}\left|\Omega\right|$
$\leq$ $2^{\frac{1}{p^{-}-1}}\left(\lambda_{2}+1\right)^{\frac{p^{-}}{p^{-}-1}}\left(M_{1}+1\right)^{p^{+}}+2^{\frac{1}{p^{-}-1}}\left|\Omega\right|$
which yields (34). Using (31), we obtain
$\left\|\nabla\sqrt{\varepsilon-\gamma}\frac{\partial u_{m}}{\partial t}\right\|^{2}_{L^{2}(Q_{T})}=$ $\sum_{j=1}^{m}\left(\varepsilon-\gamma\right)\int_{(j-1)\frac{T}{m}}^{j\frac{T}{m}}\int_{\Omega}\left|\frac{m}{T}\left(\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right)\right|^{2}dxdt$
$=$ $\sum_{j=1}^{m}\left(\varepsilon-\gamma\right)\frac{T}{m}\int_{\Omega}\left|\frac{m}{T}\left(\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right)\right|^{2}dx$
$=$ $\sum_{j=1}^{m}\left(\varepsilon-\gamma\right)\frac{m}{T}\int_{\Omega}\left|\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right|^{2}dx\leq M_{1}.$
Similarly,
$\left\|\frac{\partial u_{m}}{\partial t}\right\|_{L^{2}(Q_{T})}^{2}=\sum_{j=1}^{m}\frac{m}{T}\int_{\Omega}\left|u_{m}^{j}-u_{m}^{j-1}\right|^{2}dx\leq M_{1}.$
Hence, we arrive at inequalities (33). To prove (35), it suffices to show that
$\left\|u_{m}-\bar{u}_{m}\right\|_{L^{2}(Q_{T})}^{2}=$ $\sum_{j=1}^{m}\int_{(j-1)\frac{T}{m}}^{j\frac{T}{m}}\int_{\Omega}\left|\left(\frac{m}{T}t-j\right)\left(u_{m}^{j}-u_{m}^{j-1}\right)\right|^{2}dxdt$
$\leq$ $\sum_{j=1}^{m}\int_{(j-1)\frac{T}{m}}^{j\frac{T}{m}}\int_{\Omega}\left|u_{m}^{j}-u_{m}^{j-1}\right|^{2}dxdt$
$=$ $\left(\frac{T}{m}\right)^{2}\sum_{j=1}^{m}\frac{m}{T}\int_{\Omega}\left|u_{m}^{j}-u_{m}^{j-1}\right|^{2}dx.$
The proof is thus complete.
By Lemma 2, upon extracting suitable (not relabeled) subsequences, we obtain the following convergences as $m\to\infty$:
$u_{m}\rightharpoonup u\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;W^{1,p(\cdot)}(\Omega))$ (36)
$u_{m}\rightharpoonup u\text{ weakly in }H^{1}(0,T;H^{1}(\Omega))$
$u_{m}\rightarrow u\text{ strongly in }C([0,T];L^{2}(\Omega))$
$\sqrt{\varepsilon-\gamma}\frac{\partial u_{m}}{\partial t}\rightharpoonup\sqrt{\varepsilon-\gamma}\frac{\partial u}{\partial t}\text{ weakly in }L^{2}(0,T;H^{1}(\Omega)),$
$\bar{u}_{m}\rightharpoonup\bar{u}\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;W^{1,p(\cdot)}(\Omega)),$
$\vec{q}(x,\nabla\bar{u}_{m})\rightharpoonup\vec{\Lambda}\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;L^{p^{\prime}(\cdot)}(\Omega)).$
Since $u_{m}-\bar{u}_{m}$ converges strongly to zero in $L^{2}(Q_{T})$, we deduce that $u=\bar{u}$ a.e. in $Q_{T}$. Moreover, $u(t)\in W_{f}^{1,p(\cdot)}(\Omega)$ a.e. in $[0,T]$ follows from (36). Passing to the limit as $m\rightarrow\infty$ in (32), we then obtain
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\vec{\Lambda}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (37)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. To determine the function $\vec{\Lambda}$, we further consider the stronger convergence of the sequence $\{\nabla\bar{u}_{m}\}_{m\geq m_{0}}$, as stated in the following lemma.
Lemma 13
Let $m$ denote the index of the subsequence along which the convergences in (36) hold. Then both $\nabla u_{m}$ and $\nabla\bar{u}_{m}$ converge strongly in $L^{2}(0,T;L^{2}(\Omega))$ as $m\to\infty$.
Proof 5
We first examine the relationship between the piecewise constant and piecewise linear interpolations of a function defined at discrete time steps.
For any $i=1,2,\cdots,m$ and $t\in[(i-1)\frac{T}{m},i\frac{T}{m})$, by construction of approximate solutions, we have
$\nabla u_{m}(t)-\nabla u(t)=\nabla u_{m}^{i}-\nabla u(t)-\left(i-\frac{m}{T}t\right)\left(\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right),$
and thus
$\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u(t)\right|^{2}dx=$ $\int_{\Omega}\left|\nabla u_{m}(t)-\nabla u(t)+\left(i-\frac{m}{T}t\right)\left(\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right)\right|^{2}dx$
$\leq$ $2\int_{\Omega}\left|\nabla u_{m}(t)-\nabla u(t)\right|^{2}dx+2\left(i-\frac{m}{T}t\right)^{2}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$ (38)
Integrating the above inequality over $(i-1)\frac{T}{m},i\frac{T}{m})$ gives
$\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u\right|^{2}dxdt$
$\leq$ $2\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxdt+2\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\left(i-\frac{m}{T}t\right)^{2}dt\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx$
$=$ $2\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxdt+\frac{2T}{3m}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$ (39)
For any $t\in[0,T)$, there exists an index $l\in\{1,2,\cdots,m\}$ such that $(l-1)\frac{T}{m}\leq t<l\frac{T}{m}$. For the case $l\geq 2$, from (31), (39), and the Hölder inequality (see Lemma 3), we derive
$\int_{0}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau=$ $\sum_{i=1}^{l-1}\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u\right|^{2}dxdt+\int_{(l-1)\frac{T}{m}}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau$
$\leq$ $\sum_{i=1}^{l-1}\left[2\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxdt+\frac{2T}{3m}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx\right]$
$+\int_{(l-1)\frac{T}{m}}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau$
$\leq$ $2\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau+\frac{2T}{3m}\sum_{i=1}^{l-1}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx$
$+2\int_{(l-1)\frac{T}{m}}^{t}\left\|\nabla\bar{u}_{m}(\tau)\right\|_{L^{2}(\Omega)}^{2}d\tau+2\int_{(l-1)\frac{T}{m}}^{t}\left\|\nabla u(\tau)\right\|_{L^{2}(\Omega)}^{2}d\tau$
$\leq$ $2\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau+\frac{2T^{2}}{3m^{2}\left(\varepsilon-\gamma\right)}\sum_{i=1}^{l-1}\left(\varepsilon-\gamma\right)\frac{m}{T}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx$
$+2\int_{(l-1)\frac{T}{m}}^{t}\left(2\left(1+\left|\Omega\right|\right)\right)^{2}\left\|\nabla u_{m}(\tau)\right\|_{L^{p(\cdot)}(\Omega)}^{2}d\tau$
$+2\int_{(l-1)\frac{T}{m}}^{t}\left(2\left(1+\left|\Omega\right|\right)\right)^{2}\left\|\nabla u(\tau)\right\|_{L^{p(\cdot)}(\Omega)}^{2}d\tau$
$\leq$ $2\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau+\frac{2T^{2}}{3m^{2}\left(\varepsilon-\gamma\right)}C_{1}+\frac{T}{m}C_{1},$ (40)
where $C_{1}>0$ depends only on $\Omega$, $M_{1}$, and $M_{2}$. As for the case $l=1$, the proof details are omitted. Indeed, integrating both sides of (38) with $i=1$ over $(0,t)$ yields a result analogous to (40).
We now analyze the convergence of $\{\nabla u_{m}\}_{m\geq m_{0}}$ and $\{\nabla\bar{u}_{m}\}_{m\geq m_{0}}$ based on the preceding discussion. By choosing the test functions $\zeta=\bar{u}_{m}-u$ in (32) and $u_{m}-u$ in (37), respectively, and then subtracting the resulting identities, we have, for $t\in[0,T]$,
$\int_{0}^{t}$ $\int_{\Omega}\left(\vec{q}(x,\nabla\bar{u}_{m})\cdot\left(\nabla\bar{u}_{m}-\nabla u\right)-\vec{\Lambda}\cdot\left(\nabla u_{m}-\nabla u\right)\right)dxd\tau$
$+\varepsilon\int_{0}^{t}\int_{\Omega}\left(\nabla\frac{\partial u_{m}}{\partial\tau}\cdot\left(\nabla\bar{u}_{m}-\nabla u\right)-\nabla\frac{\partial u}{\partial\tau}\cdot\left(\nabla u_{m}-\nabla u\right)\right)dxd\tau$
$+\int_{0}^{t}\int_{\Omega}\left(\frac{\partial u_{m}}{\partial\tau}\left(\bar{u}_{m}-u\right)-\frac{\partial u}{\partial\tau}\left(u_{m}-u\right)\right)dxd\tau=0.$
We denote the three terms on the left-hand side of the above equality by $T_{1}$, $T_{2}$, and $T_{3}$, and write $\theta(m)$ for quantities tending to zero as $m\rightarrow\infty$, uniformly with respect to $t$. By (18) and Lemma 6 (e.g., with $\varPsi=\{\chi_{Q_{t}}\vec{q}(x,\nabla u):t\in[0,T]\}$), we obtain
$T_{1}=$ $\int_{0}^{t}\int_{\Omega}\left(\vec{q}(x,\nabla\bar{u}_{m})-\vec{q}(x,\nabla u)\right)\cdot\left(\nabla\bar{u}_{m}-\nabla u\right)dxd\tau-\int_{0}^{t}\int_{\Omega}\vec{\Lambda}\cdot\left(\nabla u_{m}-\nabla u\right)dxd\tau$
$+\int_{0}^{t}\int_{\Omega}\vec{q}(x,\nabla u)\cdot\left(\nabla\bar{u}_{m}-\nabla u\right)dxd\tau$
$\geq$ $-K\int_{0}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau-\sup_{0\leq t\leq T}\left|\iint_{Q_{T}}\left(\chi_{Q_{t}}\vec{\Lambda}\right)\cdot\left(\nabla u_{m}-\nabla u\right)dxd\tau\right|$
$-\sup_{0\leq t\leq T}\left|\iint_{Q_{T}}\left(\chi_{Q_{t}}\vec{q}(x,\nabla u)\right)\cdot\left(\nabla\bar{u}_{m}-\nabla u\right)dxd\tau\right|$
$\geq$ $-K\int_{0}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau-\theta(m),$
which, together with (40), implies that
$T_{1}\geq-2K\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau-\frac{\theta(m)}{\varepsilon-\gamma}-\theta(m).$
Direct calculation shows that
$T_{2}=$ $\varepsilon\int_{0}^{t}\int_{\Omega}\left(\left(\nabla\frac{\partial u_{m}}{\partial\tau}-\nabla\frac{\partial u}{\partial\tau}\right)\cdot\left(\nabla u_{m}-\nabla u\right)+\nabla\frac{\partial u_{m}}{\partial\tau}\cdot\left(\nabla\bar{u}_{m}-\nabla u_{m}\right)\right)dxd\tau$
$=$ $\varepsilon\int_{0}^{t}\int_{\Omega}\left(\nabla\frac{\partial u_{m}}{\partial\tau}-\nabla\frac{\partial u}{\partial\tau}\right)\cdot\left(\nabla u_{m}-\nabla u\right)dxd\tau+\varepsilon\int_{0}^{t}\int_{\Omega}\nabla\frac{\partial u_{m}}{\partial\tau}\cdot\left(\nabla\bar{u}_{m}-\nabla u_{m}\right)dxd\tau$
$=$ $\frac{\varepsilon}{2}\int_{\Omega}\left|\nabla u_{m}(t)-\nabla u(t)\right|^{2}dx-\frac{\varepsilon}{2}\int_{\Omega}\left|\nabla u_{m}(0)-\nabla u(0)\right|^{2}dx$
$+\varepsilon\int_{0}^{t}\int_{\Omega}\nabla\frac{\partial u_{m}}{\partial\tau}\cdot\left(\nabla\bar{u}_{m}-\nabla u_{m}\right)dxd\tau.$ (41)
It remains to estimate the last two terms on the right-hand side of (41). On the one hand, for any $t\in[0,T)$, assume that $t\in[(j-1)\frac{T}{m},j\frac{T}{m})$ for some $j=1,2,\cdots,m$. Then
$\varepsilon\left|\int_{0}^{t}\int_{\Omega}\nabla\frac{\partial u_{m}}{\partial\tau}\cdot\left(\nabla\bar{u}_{m}-\nabla u_{m}\right)dxd\tau\right|$
$\leq$ $\varepsilon\int_{0}^{j\frac{T}{m}}\int_{\Omega}\left|\nabla\frac{\partial u_{m}}{\partial t}\cdot\left(\nabla\bar{u}_{m}-\nabla u_{m}\right)\right|dxdt$
$=$ $\varepsilon\sum_{i=1}^{j}\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\frac{m}{T}\left(i-\frac{m}{T}t\right)\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dxdt$
$=$ $\varepsilon\sum_{i=1}^{j}\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\frac{m}{T}\left(i-\frac{m}{T}t\right)dt\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx$
$=$ $\frac{T\varepsilon}{2m\left(\varepsilon-\gamma\right)}\sum_{i=1}^{j}\left(\varepsilon-\gamma\right)\frac{m}{T}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx$
$\leq$ $\frac{TM_{1}\varepsilon}{2m\left(\varepsilon-\gamma\right)}.$ (42)
On the other hand, we choose the test function $\zeta=\xi\eta$ in equation (32), where $\xi\in C_{c}^{1}(\Omega)$ and $\eta\in C^{1}([0,T])$ satisfies $\eta(T)=0$ and $\eta(0)\neq 0$. Using integration by parts, we get
$\int_{\Omega}f\xi\eta(0)dx-\iint_{Q_{T}}u_{m}\xi\eta_{t}dxdt+\iint_{Q_{T}}\left(\vec{q}(x,\nabla\bar{u}_{m})\cdot\nabla\xi\eta+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\xi\eta\right)dxdt=0.$
Letting $m\rightarrow\infty$, we have from (36) that
$\int_{\Omega}f\xi\eta(0)dx-\iint_{Q_{T}}u\xi\eta_{t}dxdt+\iint_{Q_{T}}\left(\vec{\Lambda}\cdot\nabla\xi\eta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\xi\eta\right)dxdt=0.$
Similarly, substituting the test function $\zeta$ into (37), we obtain
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\xi\eta+\vec{\Lambda}\cdot\nabla\xi\eta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\xi\eta\right)dxdt=0.$
Then
$\int_{\Omega}u(0)\xi\eta(0)dx-\iint_{Q_{T}}u\xi\eta_{t}dxdt+\iint_{Q_{T}}\left(\vec{\Lambda}\cdot\nabla\xi\eta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\xi\eta\right)dxdt=0,$
which yields $u(0)=f$ a.e. in $\Omega$. In summary, combining (41) and (42), we deduce
$T_{2}\geq\frac{\varepsilon}{2}\int_{\Omega}\left|\nabla u_{m}(t)-\nabla u(t)\right|^{2}dx-\frac{\varepsilon}{\varepsilon-\gamma}\theta(m).$
For the estimate of $T_{3}$, we apply Hölder’s inequality and derive
$\left|T_{3}\right|=$ $\left|\int_{0}^{t}\int_{\Omega}\left(\left(\frac{\partial u_{m}}{\partial\tau}-\frac{\partial u}{\partial\tau}\right)\left(u_{m}-u\right)+\frac{\partial u_{m}}{\partial\tau}\left(\bar{u}_{m}-u_{m}\right)\right)dxd\tau\right|$
$\leq$ $\frac{1}{2}\int_{\Omega}\left|u_{m}(t)-u(t)\right|^{2}dx+\left|\int_{0}^{t}\int_{\Omega}\frac{\partial u_{m}}{\partial\tau}\left(\bar{u}_{m}-u_{m}\right)dxd\tau\right|$
$\leq$ $\frac{1}{2}\sup_{0\leq t\leq T}\left\|u_{m}(t)-u(t)\right\|_{L^{2}(\Omega)}^{2}+\left\|\frac{\partial u_{m}}{\partial t}\right\|_{L^{2}(Q_{T})}\left\|\bar{u}_{m}-u_{m}\right\|_{L^{2}(Q_{T})}$
$\leq$ $\frac{1}{2}\left\|u_{m}-u\right\|_{L^{\infty}(0,T;L^{2}(\Omega))}^{2}+\frac{1}{m}M_{2}^{2}.$
Hence, from (33), (35), and (36), it follows that $T_{3}=\theta(m)$. Taking into account that $T_{1}+T_{2}+T_{3}=0$ and combining this with the above estimates, we obtain
$\frac{\varepsilon}{2}\int_{\Omega}\left|\nabla u_{m}(t)-\nabla u(t)\right|^{2}dx\leq 2K\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau+\frac{\theta(m)}{\varepsilon-\gamma}+\theta(m),$
where $\theta(m)$ may be taken non-negative without loss of generality, otherwise we consider its absolute value. As a result,
$\partial_{t}\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau\leq\frac{4K}{\varepsilon}\int_{0}^{t}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxd\tau+\frac{\theta(m)}{\varepsilon\left(\varepsilon-\gamma\right)}.$
Now, applying Gronwall’s inequality, we arrive at
$\int_{0}^{T}\int_{\Omega}\left|\nabla u_{m}-\nabla u\right|^{2}dxdt\leq\frac{\theta(m)}{4K\left(\varepsilon-\gamma\right)}\exp\left(\frac{4KT}{\varepsilon}\right),$
which tends to zero as $m\to\infty$. Therefore,
$\nabla u_{m}\rightarrow\nabla u\text{ strongly in }L^{2}(0,T;L^{2}(\Omega)).$
From (40), we further deduce that
$\nabla\bar{u}_{m}\rightarrow\nabla u\text{ strongly in }L^{2}(0,T;L^{2}(\Omega)).$ (43)
The proof of Lemma 13 is complete.
Next, we proceed to establish the form of the vector-valued function $\vec{\Lambda}$.
Lemma 14
It holds that
$\vec{\Lambda}(x,t)=\vec{q}\left(x,\nabla u(x,t)\right)\text{ a.e. in }Q_{T}.$ (44)
Proof 6
Observe that for any $t\in[0,T]$, we have
$\varrho_{L^{p^{\prime}(\cdot)}(\Omega)}\left(\left|\nabla\bar{u}_{m}(t)\right|^{p(x)-2}\nabla\bar{u}_{m}(t)\right)=$ $\int_{\Omega}\left|\left|\nabla\bar{u}_{m}(t)\right|^{p(x)-2}\nabla\bar{u}_{m}(t)\right|^{\frac{p(x)}{p(x)-1}}dx$
$=$ $\int_{\Omega}\left|\nabla\bar{u}_{m}(t)\right|^{p(x)}dx$
$\leq$ $\max\left\{\left\|\bar{u}_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}^{p^{-}},\left\|\bar{u}_{m}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}^{p^{+}}\right\},$
which gives
$\sup_{0\leq t\leq T}\left\|\left|\nabla\bar{u}_{m}(t)\right|^{p(x)-2}\nabla\bar{u}_{m}(t)\right\|_{L^{p^{\prime}(\cdot)}(\Omega}\leq\left(M_{2}+1\right)^{p^{+}-1}.$
Therefore, there exists a vector-valued function $\vec{\sigma}(x,t)$ such that
$\left|\nabla\bar{u}_{m}\right|^{p(x)-2}\nabla\bar{u}_{m}\rightharpoonup\vec{\sigma}\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;L^{p^{\prime}(\cdot)}(\Omega).$
Considering the global Lipschitz continuity of the mapping $\xi\mapsto\frac{\xi}{1+\left|\xi\right|^{2}}$, we deduce from (43) that
$\frac{\nabla\bar{u}_{m}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}\rightarrow\frac{\nabla u}{1+\left|\nabla u\right|^{2}}\text{ strongly in }L^{2}(Q_{T}).$ (45)
Hence, to prove (44), it suffices to show that
$\vec{\sigma}(x,t)=\left|\nabla u(x,t)\right|^{p(x)-2}\nabla u(x,t)\text{ a.e. in }Q_{T}.$
The integral identity (32) can be rewritten as
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\zeta+\frac{\nabla\bar{u}_{m}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}\cdot\nabla\zeta+\delta\left|\nabla\bar{u}_{m}\right|^{p(x)-2}\nabla\bar{u}_{m}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (46)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Letting $m\rightarrow\infty$ in (46) and using (45), we have
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\frac{\nabla u}{1+\left|\nabla u\right|^{2}}\cdot\nabla\zeta+\delta\vec{\sigma}\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0.$ (47)
Taking $\zeta=u$, we obtain
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}u+\frac{\left|\nabla u\right|^{2}}{1+\left|\nabla u\right|^{2}}+\delta\vec{\sigma}\cdot\nabla u+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla u\right)dxdt=0.$
Furthermore, substituting $\zeta=\bar{u}_{m}$ into (46) yields
$\iint_{Q_{T}}\left(\frac{\partial u_{m}}{\partial t}\bar{u}_{m}+\frac{\left|\nabla\bar{u}_{m}\right|^{2}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}+\delta\left|\nabla\bar{u}_{m}\right|^{p(x)}+\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\bar{u}_{m}\right)dxdt=0.$ (48)
Since
$\left|\frac{\left|\nabla\bar{u}_{m}(x,t)\right|^{2}}{1+\left|\nabla\bar{u}_{m}(x,t)\right|^{2}}-\frac{\left|\nabla u(x,t)\right|^{2}}{1+\left|\nabla u(x,t)\right|^{2}}\right|\leq\left|\left|\nabla\bar{u}_{m}(x,t)\right|-\left|\nabla u(x,t)\right|\right|\leq\left|\nabla\bar{u}_{m}(x,t)-\nabla u(x,t)\right|,$
for a.e. $(x,t)\in Q_{T}$. Combining this with (43), (47), and (48), we deduce that
$\iint_{Q_{T}}\delta\vec{\sigma}\cdot\nabla udxdt=$ $-\iint_{Q_{T}}\frac{\partial u}{\partial t}udxdt-\iint_{Q_{T}}\frac{\left|\nabla u\right|^{2}}{1+\left|\nabla u\right|^{2}}dxdt-\iint_{Q_{T}}\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla udxdt$
$=$ $\lim_{m\rightarrow\infty}\left(-\iint_{Q_{T}}\frac{\partial u_{m}}{\partial t}\bar{u}_{m}dxdt-\iint_{Q_{T}}\frac{\left|\nabla\bar{u}_{m}\right|^{2}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}dxdt-\iint_{Q_{T}}\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\bar{u}_{m}dxdt\right)$
$=$ $\limsup_{m\rightarrow\infty}\iint_{Q_{T}}\delta\left|\nabla\bar{u}_{m}\right|^{p(x)}dxdt.$
Given that the mapping $\xi\mapsto\frac{1}{p}\left|\xi\right|^{p}$ is convex for $p\geq 1$, it follows that for any $\psi\in L^{\infty}(0,T;W^{1,p(\cdot)}(\Omega))$, we have
$\iint_{Q_{T}}\left(\left|\nabla\bar{u}_{m}\right|^{p(x)-2}\nabla\bar{u}_{m}-\left|\nabla\psi\right|^{p(x)-2}\nabla\psi\right)\cdot\left(\nabla\bar{u}_{m}-\nabla\psi\right)dxdt\geq 0.$
Passing to the limit as $m\rightarrow\infty$ in the above inequality, we derive
$\iint_{Q_{T}}\left(\vec{\sigma}-\left|\nabla\psi\right|^{p(x)-2}\nabla\psi\right)\cdot\left(\nabla u-\nabla\psi\right)dxdt\geq 0.$
Thus, applying the standard monotone method, we obtain $\vec{\sigma}=\left|\nabla u\right|^{p(x)-2}\nabla u$ a.e. in $Q_{T}$, which verifies that (44) holds.
Based on the results established earlier, we now prove that the regularized problem (21)–(23) admits a unique weak solution.
Proof of Theorem 2 1
Combining (37) and (44) yields the following integral identity
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\vec{q}(x,\nabla u)\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (49)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. The Neumann boundary condition of problem (21)–(23) is implicitly incorporated in the weak variational formulation $J^{m,j}$ and can be recovered via integration by parts. Next, we address the uniqueness of the weak solution to problem (21)–(23). To this end, assume that $u_{1}$ and $u_{2}$ are two weak solutions with the same initial data $f$. For any $t\in[0,T]$, choosing the test function $\zeta=\left(u_{1}-u_{2}\right)\chi_{[0,t]}(\tau)$ in (49), we obtain
$\iint_{Q_{t}}\left(\frac{\partial u_{1}}{\partial\tau}-\frac{\partial u_{2}}{\partial\tau}\right)\left(u_{1}-u_{2}\right)dxd\tau$
$+$ $\iint_{Q_{t}}\left(\frac{\nabla u_{1}}{1+\left|\nabla u_{1}\right|^{2}}-\frac{\nabla u_{2}}{1+\left|\nabla u_{2}\right|^{2}}\right)\cdot\left(\nabla u_{1}-\nabla u_{2}\right)dxd\tau$
$+$ $\delta\iint_{Q_{t}}\left(\left|\nabla u_{1}\right|^{p(x)-2}\nabla u_{1}-\left|\nabla u_{2}\right|^{p(x)-2}\nabla u_{2}\right)\cdot\left(\nabla u_{1}-\nabla u_{2}\right)dxd\tau$
$+$ $\varepsilon\iint_{Q_{t}}\left(\nabla\frac{\partial u_{1}}{\partial\tau}-\nabla\frac{\partial u_{2}}{\partial\tau}\right)\cdot\left(\nabla u_{1}-\nabla u_{2}\right)dxd\tau=0.$
The global Lipschitz continuity of the mapping $\xi\mapsto\frac{\xi}{1+\left|\xi\right|^{2}}$ implies that
$\frac{1}{2}\int_{\Omega}\left|u_{1}(t)-u_{2}(t)\right|^{2}dx+\frac{\varepsilon}{2}\int_{\Omega}\left|\nabla u_{1}(t)-\nabla u_{2}(t)\right|^{2}dx$
$\leq$ $\iint_{Q_{t}}\left(\frac{\nabla u_{1}}{1+\left|\nabla u_{1}\right|^{2}}-\frac{\nabla u_{2}}{1+\left|\nabla u_{2}\right|^{2}}\right)\cdot\left(\nabla u_{2}-\nabla u_{1}\right)dxd\tau$
$\leq$ $K(N)\int_{0}^{t}\int_{\Omega}\left|\nabla u_{1}-\nabla u_{2}\right|^{2}dxd\tau.$
Using Gronwall’s inequality, we derive that $u_{1}=u_{2}$ a.e. in $Q_{T}$, which shows the uniqueness of the weak solution to problem (21)–(23).
To prove Theorem 2, it remains to establish the energy estimate (25). By Lemma 11, for any $\gamma\in(0,\varepsilon)$, there exists $m_{0}(\gamma)>0$ such that, for all $m\geq m_{0}$ and for a.e. $t\in[0,T)$, we derive from (28) the following inequalities, where $(j-1)\frac{T}{m}\leq t<j\frac{T}{m}$ holds for some $j$:
$\int_{\Omega}\varphi(x,\nabla\bar{u}_{m}(t))dx+\int_{0}^{t}\int_{\Omega}\left|\frac{\partial u_{m}}{\partial\tau}\right|^{2}dxd\tau+\int_{0}^{t}\int_{\Omega}\left(\varepsilon-\gamma\right)\left|\nabla\frac{\partial u_{m}}{\partial\tau}\right|^{2}dxd\tau$
$\leq$ $\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx+\sum_{i=1}^{j}\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\frac{\partial u_{m}}{\partial t}\right|^{2}dxdt+\sum_{i=1}^{j}\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left(\varepsilon-\gamma\right)\left|\nabla\frac{\partial u_{m}}{\partial t}\right|^{2}dxdt$
$=$ $\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx+\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx+\sum_{i=1}^{j}\int_{\Omega}\left(\varepsilon-\gamma\right)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx$
$\leq$ $\int_{\Omega}\varphi(x,\nabla f)dx.$
Since $\frac{\partial u_{m}}{\partial\tau}\rightharpoonup\frac{\partial u}{\partial\tau}$ weakly in $L^{2}(0,t;H^{1}(\Omega))$, it follows from the weak lower semicontinuity of the $L^{2}(Q_{t})$-norm that
$\liminf_{m\rightarrow\infty}\int_{\Omega}\varphi(x,\nabla\bar{u}_{m}(t))dx+\int_{0}^{t}\int_{\Omega}\left|\frac{\partial u}{\partial\tau}\right|^{2}dxd\tau+\int_{0}^{t}\int_{\Omega}\left(\varepsilon-\gamma\right)\left|\nabla\frac{\partial u}{\partial\tau}\right|^{2}dxd\tau\leq\int_{\Omega}\varphi(x,\nabla f)dx.$
From Lemma 13, we know that $\nabla\bar{u}_{m}(t)\rightarrow\nabla u(t)$ a.e. in $\Omega$. Then, Fatou’s Lemma implies that
$\int_{\Omega}\varphi(x,\nabla u(t))dx+\int_{0}^{t}\int_{\Omega}\left|\frac{\partial u}{\partial\tau}\right|^{2}dxd\tau+\int_{0}^{t}\int_{\Omega}\left(\varepsilon-\gamma\right)\left|\nabla\frac{\partial u}{\partial\tau}\right|^{2}dxd\tau\leq\int_{\Omega}\varphi(x,\nabla f)dx.$
Given that $\gamma>0$ is arbitrary, we obtain
$\int_{\Omega}\varphi(x,\nabla u(t))dx+\int_{0}^{t}\int_{\Omega}\left|\frac{\partial u}{\partial\tau}\right|^{2}dxd\tau+\int_{0}^{t}\int_{\Omega}\varepsilon\left|\nabla\frac{\partial u}{\partial\tau}\right|^{2}dxd\tau\leq\int_{\Omega}\varphi(x,\nabla f)dx,$
for a.e. $t\in[0,T]$. This completes the proof of Theorem 2.
4 Proof of Theorem 1
In this section, we prove the existence of Young measure solutions to problem (1)–(3) using Theorem 2 and the corresponding limit process $\varepsilon\to 0^{+}$.
For any $\varepsilon\in(0,1)$, it follows from Theorem 2 that the regularized problem (21)–(23) admits a unique weak solution $u^{\varepsilon}\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$, which satisfies
$\iint_{Q_{T}}\left(\frac{\partial u^{\varepsilon}}{\partial t}\zeta+\vec{q}(x,\nabla u^{\varepsilon})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u^{\varepsilon}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (50)
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$, together with the energy estimate
$\int_{\Omega}\varphi(x,\nabla u^{\varepsilon}(t))dx+\left\|\frac{\partial u^{\varepsilon}}{\partial\tau}\right\|^{2}_{L^{2}(Q_{t})}+\left\|\nabla\sqrt{\varepsilon}\frac{\partial u^{\varepsilon}}{\partial\tau}\right\|^{2}_{L^{2}(Q_{t})}\leq\int_{\Omega}\varphi(x,\nabla f)dx,$
for a.e. $t\in[0,T]$. Then, from (17), we derive
$\varrho_{L^{p(\cdot)}(\Omega)}\left(\nabla u^{\varepsilon}(t)\right)\leq\lambda_{1}^{-1}\left(\int_{\Omega}\varphi(x,\nabla u^{\varepsilon}(t))dx+\left|\Omega\right|\right)\leq\lambda_{1}^{-1}\left(\int_{\Omega}\varphi(x,\nabla f)dx+\left|\Omega\right|\right),$
and furthermore,
$\varrho_{L^{p^{\prime}(\cdot)}(\Omega)}\left(\vec{q}(x,\nabla u^{\varepsilon}(t))\right)=$ $\int_{\Omega}\left|\vec{q}(x,\nabla u^{\varepsilon}(t))\right|^{\frac{p(x)}{p(x)-1}}dx$
$\leq$ $\int_{\Omega}2^{\frac{1}{p(x)-1}}\left(\lambda_{2}^{\frac{p(x)}{p(x)-1}}\left|\nabla u^{\varepsilon}(t)\right|^{p(x)}+1\right)dx$
$\leq$ $2^{\frac{1}{p^{-}-1}}\left(\lambda_{2}+1\right)^{\frac{p^{-}}{p^{-}-1}}\varrho_{L^{p(\cdot)}(\Omega)}\left(\nabla u^{\varepsilon}(t)\right)+2^{\frac{1}{p^{-}-1}}\left|\Omega\right|$
$\leq$ $2^{\frac{1}{p^{-}-1}}\left(\lambda_{2}+1\right)^{\frac{p^{-}}{p^{-}-1}}\lambda_{1}^{-1}\left(\int_{\Omega}\varphi(x,\nabla f)dx+\left|\Omega\right|\right)+2^{\frac{1}{p^{-}-1}}\left|\Omega\right|.$
By Lemmas 1 and 5, along with the mean invariance of $u^{\varepsilon}(t)$, we deduce that
$\left\|u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\left\|\nabla\sqrt{\varepsilon}u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\sup_{0\leq t\leq T}\left\|u^{\varepsilon}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}+\sup_{0\leq t\leq T}\left\|\vec{q}\left(x,\nabla u^{\varepsilon}(t)\right)\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}\leq M,$ (51)
where $M>0$ is independent of $\varepsilon$. Therefore, there exist a subsequence of $\{u^{\varepsilon}\}_{0<\varepsilon<1}$, still denoted by itself, and a function $u\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;L^{2}(\Omega))$ such that the following weak convergence results hold as $\varepsilon\to 0^{+}$:
$u^{\varepsilon}\rightharpoonup u\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;W^{1,p(\cdot)}(\Omega)),$
$u^{\varepsilon}\rightarrow u\text{ strongly in }C([0,T];L^{2}(\Omega)),$
$u^{\varepsilon}_{t}\rightharpoonup u_{t}\text{ weakly in }L^{2}(Q_{T}).$
For a.e. $t\in[0,T]$, let $\nu_{\cdot,t}=(\nu_{x,t})_{x\in\Omega}$ denote the Young measure generated by the sequence $\{\nabla u^{\varepsilon}(t)\}_{0<\varepsilon<1}$. Since the generating sequence satisfies the boundedness condition (14), it follows from Lemmas 8 and 9 that $\left\|\nu_{x,t}\right\|_{\mathcal{M}(\mathbb{R}^{N})}=1$ a.e. in $\Omega$. We now verify that the pair $\left(u,\nu\right)$ is a Young measure solution to problem (1)–(3). Combining Remark 1 with (51), we infer that
$\vec{q}(\cdot,\nabla u^{\varepsilon}(t))\rightharpoonup\langle\nu_{x,t},\vec{q}(x,\cdot)\rangle=\int_{\mathbb{R}^{N}}\vec{q}(\cdot,\xi)d\nu_{\cdot,t}(\xi)\text{ weakly in }L^{p^{\prime}(\cdot)}(\Omega),$
which yields
$\int_{\Omega}\vec{q}(x,\nabla u^{\varepsilon}(t))\cdot\vec{\psi}(t)dx\rightarrow\int_{\Omega}\langle\nu_{x,t},\vec{q}(x,\cdot)\rangle\cdot\vec{\psi}(t)dx\text{ a.e. in }[0,T],$
for every $\vec{\psi}\in L^{1}(0,T;L^{p(\cdot)}(\Omega))$. By the dominated convergence theorem, we obtain
$\int_{0}^{T}\int_{\Omega}\vec{q}(x,\nabla u^{\varepsilon})\cdot\vec{\psi}dxdt\rightarrow\int_{0}^{T}\int_{\Omega}\langle\nu_{x,t},\vec{q}(x,\cdot)\rangle\cdot\vec{\psi}dxdt,$
and thus
$\vec{q}(\cdot,\nabla u^{\varepsilon})\rightharpoonup\langle\nu,\vec{q}(x,\cdot)\rangle\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;L^{p^{\prime}(\cdot)}(\Omega)).$
In a similar manner, we have
$\nabla u^{\varepsilon}\rightharpoonup\langle\nu,\mathrm{id}\rangle\text{ weakly }{\ast}\text{ in }L^{\infty}(0,T;L^{p(\cdot)}(\Omega)).$
Then, we obtain $\nabla u(x,t)=\langle\nu_{x,t},\mathrm{id}\rangle$ a.e. in $Q_{T}$, proving (19). Furthermore, using Hölder’s inequality, we derive
$\left|\iint_{Q_{T}}\varepsilon\nabla\frac{\partial u^{\varepsilon}}{\partial t}\cdot\nabla\zeta dxdt\right|\leq\sqrt{\varepsilon}\left\|\nabla\sqrt{\varepsilon}\frac{\partial u^{\varepsilon}}{\partial t}\right\|_{L^{2}(Q_{T})}\left\|\nabla\zeta\right\|_{L^{2}(Q_{T})}\leq\sqrt{\varepsilon M}\left\|\nabla\zeta\right\|_{L^{2}(Q_{T})}.$
Taking the limit $\varepsilon\rightarrow 0^{+}$ in (50), we arrive at
$\iint_{Q_{T}}\left(\frac{\partial u}{\partial t}\zeta+\langle\nu_{x,t},\vec{q}(x,\cdot)\rangle\cdot\nabla\zeta\right)dxdt=0,$
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Moreover, $u(0)=f$ a.e. in $\Omega$ follows from the same argument as in the proof of Lemma 13, and the details are omitted here. Therefore, $\left(u,\nu\right)$ is a Young measure solution to problem (1)–(3), and the proof is complete.
Acknowledgements
The authors wish to express their sincere gratitude to Lecturer Jingfeng Shao (Guangxi University) for many helpful discussions during the preparation of this paper. This work was supported by the National Natural Science Foundation of China (Grant Nos. U21B2075, 12171123, 12301536) and the Fundamental Research Funds for the Central Universities (Grant Nos. 2022FRFK060020, HIT.NSRIF202302).
Data availability
No data was used for the research described in the article.